OpenStudy (anonymous):

Which property of addition is shown below? If x = a + bi and y = –a – bi, x + y = 0. a. commutative property b. identity property c. associative property d. inverse property Thinking a?

OpenStudy (nikato):

I don't think so.a would just be something like a+b=b+a Should be inverse, D. a+-a=0 and bi-bi=0. -a is the inverse of a and -bi is the inverse of bi